OpenStudy (anonymous):

WHAT IS THE SOLUTION to 2x^2+x+2=0?

OpenStudy (anonymous):

by adding them both together.

OpenStudy (anonymous):

umm, hello?

OpenStudy (anonymous):

it has imaginary roots

OpenStudy (anonymous):

\[x=-0.25\pm 0.968i\]

OpenStudy (phi):

For this problem, you should use the quadratic formula see http://www.khanacademy.org/math/trigonometry/polynomial_and_rational/quad_formula_tutorial/v/using-the-quadratic-formula for how to do it. You will get a minus number inside a square root, which means a complex number.

OpenStudy (anonymous):

isn't the answer x=-1+isqrt15/2?

OpenStudy (phi):

almost, the bottom should be 2*a = 4 \[ x= \frac{-1 ± \sqrt{15} \ i}{4} \]